OpenStudy (anonymous):

what is the area of right triangle whose sides haves lenghts 10,24,26

OpenStudy (shamim):

Area of a tringle=0.5*base*height=0.5*10*24=?

OpenStudy (anonymous):

area triangle 1/2 base times height 1/2*24* 10 =120

OpenStudy (michele_laino):

please use the formula of Erone

OpenStudy (anonymous):

Heron's Formula \[area = \sqrt{s(s-a)(s-b)(s-c) }\]here s = semi-perimeter \[s = \frac{a+b+c}{2}\] for side lengths a, b, c.

OpenStudy (michele_laino):

hint: \[\begin{gathered} S = \sqrt {p\left( {p - a} \right)\left( {p - b} \right)\left( {p - c} \right)} \hfill \\ {\text{where }}p{\text{ is the half - perimeter}}{\text{, and}} \hfill \\ a,\;b,\;c{\text{ are the lengths of the sides of the triangle}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

the triangle is a right triangle @shamim caught it first. The smaller sides have to be perpendicular to each other. So, Area = (1/ 2) * base * height = (1/2)*10*24 = 120 sq. units

OpenStudy (michele_laino):

you are right! @akitav